OpenStudy (anonymous):

Plaese solve exponential equation: 4 over 9 exponent of C+3 = 3 over 2 exponent of C+9

OpenStudy (anonymous):

How about writing that in a mathematical notation?

OpenStudy (anonymous):

\[\left(\begin{matrix}4 \over\ 9\end{matrix}\right)^C+^3 = \left(\begin{matrix}3 \over\ 2\end{matrix}\right)^C+^9\]

OpenStudy (anonymous):

c = -5

OpenStudy (anonymous):

WHAT!! How did u get that??

OpenStudy (anonymous):

Can you explain please.

OpenStudy (anonymous):

Do you know logs?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That's aweome. \[\log(4/9)^{C+3} = \log(3/2)^{C+9}\]

OpenStudy (anonymous):

Then, \[(C+3) \log(4/9) = (C+9) \log (3/2)\]

OpenStudy (anonymous):

Then, \[(C+3)/(C+9) = \log(3/2) / \log(4/9)\]

OpenStudy (anonymous):

That would make it equal -7

OpenStudy (anonymous):

\[(C+3)/(C+9)=-0.5\]

OpenStudy (anonymous):

oh k i got it thank you.

OpenStudy (anonymous):

Cool, solve for C.